Question 3) 1 point Calculate elasticity when P 20; and Qd 400-10P & Write its interpretation
Answer #1

Ans. P=20, Q d= 400-10p

quantity demand (Qd), when P=20

Qd= 400 - 10( 20)= 400 - 200 = 200

Qd= 400-10p

Differentiation w.r.t. 'P', we have

dQ_ -10 dp 2

The elasticity of demand (Ed ) = d2

= - 10 X 20/ 200

= -1

hence, the elasticity of demand is unitary elastic. which means that when price increases by 1 %, quantity demanded decreases by 1% (a negative relationship between quantity demand and price of a good) and vice versa.